a circular dartboard has a radius of 2 meters and a red circle in the center. Assume you hit the target at a ramdom point. For what radius of the center regions does P(hitting red)= 0.6?
Answer by stanbon(75887)

a circular dartboard has a radius of 2 meters and a red circle in the center. Assume you hit the target at a ramdom point. For what radius of the center regions does P(hitting red)= 0.6?
----------------------
Solve:
Area of red/total area = 0.6
----
Area of red = 0.6(pi*2^2)
----
Area of red = 0.6(4pi) = 2.4pi sq in.
----
Solve: (pi)r^2 = 2.4(pi)
r^2 = 2.4
r = 1.55 inches
======================
Cheers,
Stan H.
======================